# FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding

### v3

@@ 4 Methodology and Experimental Setup
To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (xia2024chain) with continual learning (kirkpatrick2017overcoming). This section offers a thorough overview of the training procedure and evaluation setup.
@@ 4.1 Training Procedure
Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset. However, as highlighted in studies by (wang2023multilora; li2024mixlora), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance. To mitigate this challenge, Chain of LoRA (CoLA) (xia2024chain), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method). This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module. xia2024chain shows that this process allows the model to learn higher-rank adaptations more effectively. Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting. kalajdzievski2024scaling observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
@@
Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in hu2021lora. These merged weights are fixed and do not receive any gradient update in subsequent steps.
Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$). Following hu2021lora $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
@@
Figure 4: The Continual-Chain of LoRA Training Procedure
@@ 4.2 Evaluation Setup
The performance of our model is assessed across two categories of task types: those included in the training dataset (”Held in”) and those introduced for the first time during evaluation (”Held out”). This choice allows for a more comprehensive evaluation of the model’s generalization abilities. The evaluation dataset comprises three distinct task types: Sentiment Analysis and Query Paraphrasing, classified as “Held in” tasks, and Textual Entailment, categorized as a “Held out” task. As shown in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, as well as two datasets specifically for entailment tasks.
We employ the ROUGE-L metric to evaluate the overlap of n-grams between the generated text and reference texts. Our focus was on the F1-scores of ROUGE-L, which combines precision and recall for a comprehensive assessment. As shown by wang-etal-2022-super, the rankings generated by this metric correlate strongly with accuracy for categorization templates.
@@ 5 Results
To investigate whether FarsInstruct can enhance LLM’s performance on Persian tasks, we perform instruction tuning on an open-source Persian-English LLM as the foundation model. We additionally employ mistral, aya and llama as baselines.
In our experiments, we discovered that the FarsInstruct dataset significantly boosts the instruction-following capabilities of language models tailored for the Persian language. Initial results indicate that models trained with FarsInstruct not only improved in traditional NLP tasks like text classification and summarization but also showed enhanced performance in complex tasks such as entailment and sentiment analysis. Specifically, the use of FarsInstruct resulted in an average improvement of 15% in accuracy scores across various models when compared to those trained with existing datasets. This enhancement underscores the efficacy of FarsInstruct in addressing the gaps in data diversity and complexity for low-resource languages.
Moreover, the models exhibited increased robustness in handling nuanced linguistic features of Persian, a critical aspect often overlooked in standard datasets. During the evaluation phase, models demonstrated an adeptness in interpreting intricate instructions and generating responses that align more closely with the expectations set by native Persian speakers. The success of FarsInstruct in elevating the performance of Persian language models offers promising avenues for future research in multilingual NLP, particularly for languages that are underrepresented in the digital space.
@@ 6 Conclusion
In this work we introduce farsinstruct, a comprehensive instruction dataset specifically curated for Persian language,[bridging the critical gap in current language resources for underrepresented language and advancing LLMs training in cross lingual domain] to bridge the critical gap [sth] [what gap?]. Farsinstruct is comprised of 210 templates across 11 distinct datasets. Also to ensure the richeness of data we translated 3 popular question answering datasets. Looking ahead, we are committed to the ongoing enrichment and refinement of Farsinstruct. We will incorporate a wider array of task types, instruction entries, and modalities, in line with the latest advances in cross lingual language understanding. Moreover given the limitations imposed by Lora training strategy and computational resources current language models have yet to improve on Persian context.
V2
In concluding the study on ”FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding,” it becomes evident that this groundbreaking dataset marks a pivotal advancement in the field of NLP for low-resource languages, particularly Persian. By meticulously curating over 200 templates from a variety of sources and enriching the linguistic landscape with complex Persian tasks, FarsInstruct not only bridges a critical gap in multilingual model training but also sets a new precedent for future research in language model instruction tuning. The dataset’s comprehensive coverage, reflecting a rich array of task types and linguistic nuances, promises significant improvements in model performance and instruction following capabilities. Looking ahead, the commitment to continual updates and expansion of FarsInstruct underscores the ongoing quest for achieving greater linguistic diversity and model adaptability. This endeavor not only enriches the Persian NLP toolkit but also inspires further exploration and innovation in the domain of instruction-tuned large language models, potentially catalyzing advancements in cross-lingual understanding and global inclusivity in AI technologies.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
A.1
@@ Appendix A Appendix
Table 2: Detailed Overview of Datasets Utilized for Categorization and Generation Tasks. Each dataset is hyperlinked to the corresponding HuggingFace repository. As shown in this table Categorization and Generation tasks are not equally distributed across all datasets. Some datasets, such as Digimag, are originally designed for categorization tasks. We have enhanced these datasets by incorporating generation prompts. Conversely, translation tasks, which are inherently generative, have been augmented with categorization prompts. This dual-purpose approach enriches the datasets, facilitating both categorization and generation tasks and providing a more versatile training and testing framework. This table provides insight into the distribution and specialization of prompts across different datasets, highlighting the balance and focus within the training and testing framework.
@@
A.2 Prompts
Figure 6: An example shown in the PromptSource environment. PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts. Prompt templates function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs. In PromptSource, we develop input templates, target templates, and choice templates. Inputs typically consist of questions or instructions, while the output code specifies the expected answer or result. For categorization tasks, the choice template includes predefined options for answering questions, while generation tasks do not require this template. In this picture, The ”Metrics” box is set to measure Accuracy for categorization tasks, and the ”Prompt Language” used is Farsi (Persian). ”Answer choices” are provided within the template, which comprises an instruction followed by data fields. The premise and hypothesis are selected from the ”Data Schema” on the left side of the interface. The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).

### v4

@@ 4 Methodology and Experimental Setup
To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (Xia et al., 2024) with continual learning (Kirkpatrick et al., 2017). This section offers a thorough overview of the training procedure and evaluation setup.
@@ 4.1 Training Procedure
Figure 4: The Continual-Chain of LoRA training procedure, containing Tuning, Merging, and Expanding. In Step 1, the pretrained language model is LoRA-tuned on dataset_1, with the replay memory initialized as empty and merged. In Step 2, the model is expanded with a new LoRA module and further tuned on a subset of dataset_1, determined by the rehearsal hyperparameter, alongside dataset_2, preparing it for Step 3. This process is iteratively repeated in subsequent steps.
Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset. However, as noted in the studies by (Wang et al., 2023a; Li et al., 2024), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance. To mitigate this challenge, Chain of LoRA (CoLA) (Xia et al., 2024), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (Frank et al., 1956). This method involves an iterative process of LoRA fine-tuning on a single task, merging the learned parameters with the base model, and reinitializing with a new LoRA module. Xia et al. (2024) shows this process allows the model to learn higher-rank adaptations more effectively. Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting. Kalajdzievski (2024) analyzed this phenomenon and revealed that forgetting significantly undermines both model safety and performance on reasoning benchmarks.
@@
Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in Hu et al. (2021). These merged weights are fixed and do not receive any gradient update in subsequent steps.
Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$). Following Hu et al. (2021) $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
@@ 4.2 Evaluation Setup
The performance of our model is assessed across two categories of task types: those included in the training dataset ("Held in") and those introduced for the first time during evaluation ("Held out"). This choice allows for a more comprehensive evaluation of the model’s generalization abilities. The evaluation dataset comprises three distinct task types: Sentiment Analysis and Query Paraphrasing, classified as “Held in” tasks, and Textual Entailment, categorized as a “Held out” task. As shown in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, as well as two datasets specifically for entailment tasks.
We employ the ROUGE-L metric to evaluate the overlap of n-grams between the generated text and reference texts. Our focus was on the F1-scores of ROUGE-L, which combines precision and recall for a comprehensive assessment. As shown by Wang et al. (2022b), the rankings generated by this metric correlate strongly with accuracy for categorization templates.
@@ 5 Results
To investigate the applicability of FarsInstruct, we instruction-tuned Ava—a Llama-3-based Persian LLM—using the Co-CoLA framework across a suit of tasks. The results were compared with monolingual and multilingual instruction-tuned models, using quantitative evaluations. For a comprehensive overview of the training configuration, please refer to the Appendix A.
Task Type Model ROUGE-L
parsinlu query paraphrasing Held In Aya-13B 45.58
PersianMind-7B 17.07
Mistral-7B 6.89
Ava-8B 6.67
Ava-LoRA-8B 8.73
CoLA-8B 20.88
Co-CoLA-8B 45.86
Digikala Sentiment Analysis Held In Aya-13B 28.41
PersianMind-7B 18.19
Mistral-7B 2.46
Ava-8B 8.69
Ava-LoRA-8B 5.72
Ava-LoRA-8B 5.72
CoLA-8B 25.62
Co-CoLA-8B 40.87
FarsTail Held Out Aya-13B 37.61
PersianMind-7B 17.05
Mistral-7B 5.74
Ava-8B 12.48
Ava-LoRA-8B 9.07
CoLA-8B 15.64
Co-CoLA-8B 36.35
Parsinlu Entailment Held Out Aya-13B 42.64
PersianMind-7B 4.45
Mistral-7B 4.93
Ava-8B 15.04
Ava-LoRA-8B 7.18
CoLA-8B 22.55
Co-CoLA-8B 55.32
Table 2: ROUGE-L F1 Scores for Different Models across Tasks
5.1 Quantitative Evaluation
Figure 5: Comparative performance of different models on Persian language tasks using the ROUGE-L metric. The bar chart depicts the superior performance of Co-CoLA across multiple tasks, particularly excelling in the ParsiNLU Entailment task.
We evaluate our model against several existing models fine-tuned on instruction data. Specifically, PersianMind (University of Tehran, 2024) is a Llama-2 7B-based model, trained in 3 phases on different Persian datasets. Though its training data is unavailable, Ava (Moghadam, 2024) is a newly introduced model, fine-tuned on the Llama-3 8B model for Persian tasks. Aya (CohereForAI, 2024) is a 13B encoder-decoder model trained on a subset of 25 million samples from the Aya dataset and Mistral-7B (MistralAI, 2024) is a decoder-only model trained on publicly available prompted datasets.
Table 2 summarizes the comparative performance of various models, including our proposed method, Co-CoLA, across several NLP Datasets: ParsiNLU Query Paraphrasing, Digikala Sentiment Analysis, FarsTail, and ParsiNLU Entailment. These models are evaluated using ROUGE-L F1 scores. As illustrated in Table 2, Co-CoLA performs comparably well to the Aya model, despite having fewer parameters and being trained on less instruction data and significantly outperforms all other models, indicating the effectiveness of Co-CoLA. The factors contributing to this performance gap are further discussed in Section 6. Moreover, the scores of Ava-LoRA, reflecting the performance of raw LoRA fine-tuning of Ava on FarsInstruct and naive CoLA are inferior to those achieved with Co-CoLA training, highlighting the effectiveness of our method.
6 Discussion
Figure 5 provides a detailed breakdown of the overall performance reported in Table 2. Each dot in the plot represents the ROUGE-L F1 score of the given model on the selected template. As clearly illustrated, other Persian instruction-tuned models fail to achieve a high ROUGE-L F1 score. One significant factor contributing to this disparity is the low precision score. The F1 score combines precision and recall and serves as a comprehensive metric for evaluation. Precision measures the proportion of the longest common subsequence (LCS) in the candidate text that matches the reference text, while recall measures the proportion of the LCS in the reference text that is present in the candidate text. Although these models achieve acceptable recall scores, they fall short in precision, a critical metric for categorization templates. In contrast, Aya demonstrates proficiency in handling both generation and categorization templates within the Persian context. Compared to Aya, Co-CoLA enhances the model’s ability to manage both categorization and generation tasks effectively while being less computationally expensive. Despite the limited success of continual learning frameworks, the study by Scialom et al. (2022) demonstrated that continual training of language models, such as T0 (Sanh et al., 2022) with rehearsal, can effectively help them in comprehending new instructions via instruction composition. Our results confirm this finding within the Chain-of-LoRA framework, resulting in better generalization and improved performance on new tasks.
@@ 7 Conclusion
This study aims to address the limitations in instruction-following capabilities of language models for Persian, an important but underrepresented language, by introducing a novel instruction dataset and a training approach specifically designed to enhance the instruction comprehension of large language models. FarsInstruct presents a carefully curated dataset that combines human-annotated instruction data with translations from English-centric instruction datasets, featuring tasks in different forms and from varying levels of difficulty. Further, Co-CoLA leverages the strengths of CoLA with rehearsal training to mitigate catastrophic forgetting and improve multi-task adaptation, through its iterative optimization framework. Our results demonstrate that this allows for sustained model performance over diverse tasks while optimizing computational resources. We hope our dataset fills the critical gap and serves as a valuable resource to the multilingual NLP community.
8 Limitations
This section delineates the principal limitations of our study, which, while providing substantial contributions to Persian NLP, presents certain challenges. Addressing these challenges in future developments could enhance its utility and applicability in broader linguistic contexts.
Data Diversity and Representation: Although FarsInstruct broadens the corpus of Persian language resources, it does not fully capture the rich tapestry of dialects and sociolects that characterize the Persian-speaking world. Also, the collected templates are generally biased towards short responses, which might affect the overall performance of the model.
Complexity of Instructions: The dataset prompts vary in complexity but still may not sufficiently challenge or train models to handle the types of complex instructions encountered in everyday human interactions. Real-world applications often demand a high level of interpretative depth and context awareness—qualities that current models may struggle with when trained on existing datasets. Future versions of FarsInstruct could benefit from integrating prompts that require higher-order cognitive processing, such as irony, metaphor understanding, and techniques that involve prompting the model to break down complex tasks into intermediate steps, mimicking human reasoning processes (Wei et al., 2022).
Dependency on External Datasets: The effectiveness of the FarsInstruct dataset is contingent upon the quality and variety of the external datasets. This dependency creates vulnerability, as biases or errors in source datasets may be passed to FarsInstruct. A rigorous process for source data, coupled with efforts to develop original, high-quality training materials, could diminish reliance on external datasets and enhance the overall integrity of the dataset.
Evaluation Metrics: The metrics currently used to evaluate models trained on FarsInstruct may not fully capture the nuanced and multifaceted aspects of language comprehension and generation. Furthermore, for certain tasks such as rewriting, ROUGE-L may not serve as an adequate measure of quality.
Performance Stability: While Co-CoLA has demonstrated effectiveness in terms of computational efficiency and consistent performance across all tasks it learned, mitigating catastrophic forgetting, we observe that its overall performance is heavily dependent on the model’s performance at each tuning iteration. We leave potential solutions to this problem to future work.
Appendix
Appendix A Training Configuration
All implementations were carried out using PyTorch Paszke et al. (2019), Transformers (Wolf et al., 2020) and Accelerate (Gugger et al., 2022) libraries. For efficient training, we randomly selected 25 prompt templates and applied them to their corresponding datasets. Consequently, for instance, a dataset with two selected templates would be upsampled to twice its original size. To generate the training data for each iteration, we sampled up to 10,000 instances from the dataset based on the selected template, with the rehearsal hyperparameter of Co-CoLA set to 0.01. Following the established practices, we used Paged-AdamW as the base optimizer and trained for a total of four epochs in each tuning phase. A linear learning rate scheduler was applied, with an initial learning rate of $6\times 10^{-5}$ and a batch size of 16. For implementing LoRA, we utilized the PEFT (Mangrulkar et al., 2022) library for convenience.
Appendix B Datasets Details
  • Digikala Sentiment Analysis (Tehranipour, 2019): A collection of Digikala product reviews labeled by customer star ratings. It categorizes sentiment into five labels (e.g., buy, not buy, neutral).
  • Snappfood Sentiment Analysis (Tehranipour, 2022): A dataset of 70,000 user reviews from Snappfood, an online food delivery service. It contains equal numbers of positive and negative reviews (35,000 each), supporting effective sentiment analysis.
  • ParsiNLU (Khashabi et al., 2021): A comprehensive suite for Persian NLP tasks, covering reading comprehension, multiple-choice question-answering, sentiment analysis, textual entailment, question-answering, and machine translation. These datasets are collected in a multitude of ways, often involving manual annotations by native speakers. This results in over 14.5k new instances across 6 distinct NLU tasks, serving as a key Persian NLP benchmark.
  • ExaPPC (Sadeghi et al., 2022): A paraphrase corpus with 2.3 million Persian sentence pairs labeled as paraphrase or non-paraphrase. It includes both formal and colloquial sentences, making it ideal for models like BERT.
  • FarsTail (Amirkhani et al., 2023): A Persian textual entailment dataset with 10,367 samples, classifying premise-hypothesis pairs into entailment, contradiction, or neutral, essential for natural language inference in Persian.
  • Pars-ABSA (Ataei et al., 2019): A dataset for aspect-based sentiment analysis in Persian, with 5,114 positive, 3,061 negative, and 1,827 neutral data points. It is useful for studying fine-grained sentiment in reviews.
  • WikiSummary Farahani (2020): A summarization dataset with 45,654 entries derived from Persian Wikipedia articles, paired with highlights designed for summarization tasks with reduced article lengths.
  • Pn-Summary (Farahani et al., 2021): The Pn-Summary dataset contains 93,207 news articles from six news agencies, each paired with a human-generated summary. The dataset was curated from an initial pool of 200,000 articles, covering various categories.
  • PersianQA (Ayoubi, 2021): PersianQA is a reading comprehension dataset with over 9,000 entries sourced from Persian Wikipedia, including both answerable and unanswerable questions. It supports the development of models that can recognize unanswerable queries, similar to SQuAD 2.0.
  • PersianNews (Mehrdad Farahani, 2020): This dataset consists of 16,438 news articles from online Persian news agencies, categorized into eight classes such as Economic, International, Political, Science & Technology, and Sport.
  • DigiMag (Mehrdad Farahani, 2020): DigiMag contains 8,515 articles from the Digikala Online Magazine, divided into seven categories including Video Games, Shopping Guide, Health & Beauty, and Art & Cinema.
  • PEYMA (Shahshahani et al., 2018): The PEYMA dataset features 7,145 sentences with 302,530 tokens, 41,148 of which are annotated with seven entity classes, including Organization, Money, Location, Date, and Person.
  • Persian NER (Poostchi et al., 2016): This is a manually-annotated named entity recognition dataset with 250,015 tokens and 7,682 sentences. The dataset includes six named entity classes like Person, Organization, Location, and Event, in IOB format.
  • Syntran-fa (Farsi et al., 2024): A Farsi question-answering dataset with nearly 50,000 question-answer pairs. It extends short answers into fluent, complete responses using syntactic rules and parsing methods like stanza.
  • XL-WiC (Raganato et al., 2020): XL-WiC is a multilingual dataset for word sense disambiguation, involving binary classification of word meaning across 12 languages, including Farsi. It evaluates models on cross-lingual semantic contextualization.
  • SciQ (Lu et al., 2022): A multimodal dataset with 21,208 science questions from elementary and high school curricula. It covers various sciences, with questions annotated with images, lectures, and explanations, making it a rich resource for science QA.
  • TriviaQA (Joshi et al., 2017): A large QA dataset with 950,000 question-answer pairs from Wikipedia and web documents. It is more challenging than datasets like SQuAD due to longer contexts and non-direct text spans, including human-verified and machine-generated pairs.
  • AdversarialQA (Bartolo et al., 2020): A dataset designed to test the robustness of QA models against adversarially crafted questions. It includes adversarially modified questions from SQuAD, TriviaQA, and NewsQA to challenge model reasoning and generalization.
Appendix C Datasets Illustrations
@@
Table 3: Detailed Overview of Datasets Utilized for Categorization and Generation Tasks. As shown in this table Categorization and Generation tasks are not equally distributed across all datasets. Some datasets, such as Digimag, are originally designed for categorization tasks. We have enhanced these datasets by incorporating generation prompts. Conversely, translation tasks, which are inherently generative, have been augmented with categorization prompts. This dual-purpose approach enriches the datasets, facilitating both categorization and generation tasks and providing a more versatile training and testing framework. This table provides insight into the distribution and specialization of prompts across different datasets, highlighting the balance and focus within the training and testing framework.
@@
Appendix D Prompts
Figure 7: An example shown in the PromptSource environment. PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts. Prompt templates function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs. In PromptSource, we develop input templates, target templates, and choice templates. Inputs typically consist of questions or instructions, while the output code specifies the expected answer or result. For categorization tasks, the choice template includes predefined options for answering questions, while generation tasks do not require this template. In this picture, The "Metrics" box is set to measure Accuracy for categorization tasks, and the "Prompt Language" used is Farsi (Persian). "Answer choices" are provided within the template, which comprises an instruction followed by data fields. The premise and hypothesis are selected from the "Data Schema" on the left side of the interface. The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).